Julie, Inc., has a total debt ratio of 0.20. In addition, the company had additions to retained earnings for the year just en2.4 a. (2 points) What are the dividends per share? b. (2 points) What is book value per share? c. (3 points) If the stock cu300,000 shares outstanding
Answer #1

a

DPS = dividend/ shares = 250000/300000= 0.8333

b

BVPS = ending equity/shares = 4000000/300000 = 13.33

c

Price to book = price/BVPS = 25/13.333 = 1.875

d


P/E = price*shares/(retained earnings+dividends) = 25*300000/(400000+250000) = 11.54